# Accurate Chemistry Collection: Coupled cluster atomization energies for broad chemical space

### v2

@@ Methods
Using the fully optimized B3LYP-D3(BJ)/def2-TZVPP structures, we obtain the nonrelativistic, all-electron CCSD(T)/CBS TAE using the high-level W1-F12 composite wavefunction protocol.43 W1-F12 theory is an explicitly correlated71, 72, 73 version of the original W1 theory.42 The computational details of W1-F12 theory have been specified and rationalized in great detail elsewhere.43 In brief, the Hartree–Fock (HF) energy is extrapolated to the complete basis-set limit from the cc-pVDZ-F12 and cc-pVTZ-F12 basis sets,74 using the two-point extrapolation formula, with = 5. The complementary auxiliary basis set (CABS) singles correction is included in the Hartree–Fock (HF) energy.75, 76, 77 The valence CCSD-F12 correlation energy (denoted by CCSD hereinafter) is extrapolated from the same basis sets, using an optimal extrapolation exponent for elements up to argon of = 3.67. The valence parenthetical connected triple excitations (denoted by (T)) is obtained from standard CCSD(T) calculations, as is the case in W1w theory.42, 22 Specifically, the (T) component is extrapolated from the jul-cc-pV(D+d)Z and jul-cc-pV(T+d)Z basis sets using the above two-point extrapolation formula with = 3.22.78, 79, 80, 81 The CCSD inner-shell contribution is calculated with the core-valence weighted correlation-consistent cc-pwCVTZ basis set of Peterson and Dunning,82 while the (T) inner-shell contribution is calculated with the cc-pwCVTZ basis set without the f functions. Since we are primarily interested in assessing the performance of nonrelativistic computational chemistry methods on the electronic potential energy surface, the scalar relativistic, spin-orbit, and diagonal Born-Oppenheimer corrections are not considered in the present work.
@@
The data records for 76,879 molecules of the first public version of MSR-ACC are available in Zenodo (https://doi.org/10.5281/zenodo.15387280) under the CDLA-Permissive-2.0 license. To aid machine-learning applications,83 we release also training and validation subsets of the full dataset, which constitute canonical 99% and 1% splits, respectively, of MSR-ACC/TAE25 after removing the overlap with popular W4-17 and GMTKN55 benchmark sets. The overlap contains any atomization reaction where the reconstructed molecular graph of the molecules match.
@@
It is well established that the CCSD(T) method cannot generally achieve chemical accuracy for systems dominated by moderate-to-severe nondynamical correlation effects.22, 24, 25, 23, 9, 21, 26, 84, 85, 86, 87, 88 It is therefore essential to estimate the contributions to the TAEs from post-CCSD(T) excitations. A highly successful a priori predictor for such contributions is the so-called %TAE[(T)] diagnostic,22, 23, 9 which is defined as the percentage of the TAE accounted for by parenthetical connected triple excitations:

### v3

@@ Methods
Using the fully optimized B3LYP-D3(BJ)/def2-TZVPP structures, we obtain the nonrelativistic, all-electron CCSD(T)/CBS TAE using the high-level W1-F12 composite wavefunction protocol.43 W1-F12 theory is an explicitly correlated71, 72, 73 version of the original W1 theory.42 The computational details of W1-F12 theory have been specified and rationalized in great detail elsewhere.43 In brief, the Hartree–Fock (HF) energy is extrapolated to the complete basis-set limit from the cc-pVDZ-F12 and cc-pVTZ-F12 basis sets,74 using the two-point extrapolation formula, with = 5. The complementary auxiliary basis set (CABS) singles correction is included in the Hartree–Fock (HF) energy.75, 76, 77 The valence CCSD-F12 correlation energy (denoted by CCSD hereinafter) is extrapolated from the same basis sets, using an optimal extrapolation exponent of = 3.67 for elements up to argon. Optimal values for the geminal Slater exponents were taken from Peterson et al. 74 and Hill et al. 78, and the resolution of the identity (RI) approximation was applied using the OptRI auxiliary basis sets.79 The diagonal, fixed-amplitude 3C(FIX) ansatz and the CCSD-F12b approximation were used in these calculations.76, 77 The valence parenthetical connected triple excitations (denoted by (T)) is obtained from standard CCSD(T) calculations, as is the case in W1w theory.42, 22 Specifically, the (T) component is extrapolated from the jul-cc-pV(D+d)Z and jul-cc-pV(T+d)Z basis sets using the above two-point extrapolation formula with .80, 81, 82, 83 The CCSD inner-shell contribution is calculated with the core-valence weighted correlation-consistent cc-pwCVTZ basis set of Peterson and Dunning,84 while the (T) inner-shell contribution is calculated with the cc-pwCVTZ basis set without the f functions. Since we are primarily interested in assessing the performance of nonrelativistic computational chemistry methods on the electronic potential energy surface, the scalar relativistic, spin-orbit, and diagonal Born-Oppenheimer corrections are not considered in the present work. All the calculations involved in W1-F12 theory were carried out with the Molpro 2024.1 program suite.85, 86, 87
@@
The data records for 76,879 molecules of the first public version of MSR-ACC are available in Zenodo (https://doi.org/10.5281/zenodo.15387279) under the CDLA-Permissive-2.0 license. To aid machine-learning applications,88 we release also training and validation subsets of the full dataset, which constitute canonical 99% and 1% splits, respectively, of MSR-ACC/TAE25 after removing the overlap with popular W4-17 and GMTKN55 benchmark sets. The overlap contains any atomization reaction where the reconstructed molecular graph of the molecules match.
@@
It is well established that the CCSD(T) method cannot generally achieve chemical accuracy for systems dominated by moderate-to-severe nondynamical correlation effects.22, 24, 25, 23, 9, 21, 26, 89, 90, 91, 92, 93 It is therefore essential to estimate the contributions to the TAEs from post-CCSD(T) excitations. A highly successful a priori predictor for such contributions is the so-called %TAE[(T)] diagnostic,22, 23, 9 which is defined as the percentage of the TAE accounted for by parenthetical connected triple excitations: